Find the equation of the line that passes through (-2, -3) and whose slope is 2/5.
Answer by mananth(16946)

the equation in slope point form is
y-y1=m(x-x1)
m=2/5 (x1,y1) -(-2,-3)
y-(-3)=(2/5)(x-(-2))
y+3=(2/5)(x+2)
multiply by 5
5y+15 = 2(x+2)
5y+15=2x+4
2x-5y=11